OpenStudy (anonymous):

What is the solution to the equation 2 ^1 – 2x = 3 ?

OpenStudy (anonymous):



OpenStudy (anonymous):

use logarithams \[1-2x=\log_{2}3 \]

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

\[(1-2x) \log_{10} 2 = \log_{10} 3 \] \[1-2x = (\log_{0}3) / \log_{10} 2\] \[1-2x = 1.585\] \[1-1.585 = 2x\] x = -0.2925